# MOMENT: A Family of Open Time-series Foundation Models

### v2

@@ 1 Introduction
Figure 1: MOMENT can solve multiple time series analysis tasks well (App. E).
@@ 3.2 Model Architecture
Our transformer encoder retains the modifications proposed by Raffel et al. (2020) to the original Transformer (Vaswani et al., 2017). Specifically, we remove the additive bias from the Layer Norm (Ba et al., 2016), and place it before the residual skip connections (He et al., 2016), and use the relation positional embedding scheme (Shaw et al., 2018). Below we summarize the intuition behind some of our key design decisions.
@@
Reconstruction-based anomaly detection with window size =512=512= 512
@@ 3.3 Pre-training using Masked Time series Modeling
We use the Adam optimizer with weight decay (Loshchilov & Hutter, 2019) with $\lambda=$ 0.05, $\beta_{1}=$ 0.9, $\beta_{2}=$ 0.999. We clip the gradient at 5.0, train models using a batch size of 2048, and use cosine learning rate schedule with initial and final learning rates of $1e^{-4}$ and $1e^{-5}$, respectively. We use gradient checkpointing (Radford et al., 2021) to improve training throughput and save memory, and train all models in a mixed precision setting, using float-32 for numerically unstable operations, e.g. layer normalization, and bfloat-16444https://cloud.google.com/tpu/docs/bfloat16, otherwise. We train all models for 2 epochs.
@@ 3.4 Fine-tuning on Downstream Tasks
MOMENT can be seamlessly used for multiple time series analysis tasks. In this work, we consider 5 practical time series analysis tasks as examples, namely: long- and short-horizon forecasting, classification, anomaly detection, and imputation. For forecasting tasks with horizon $H$, we replace the reconstruction head with a forecasting head, which first flattens all the $N$ $D$-dimensional patch embeddings into a $N\times D$ dimensional vector, and then projects it into a $H$-dimensional time series via a linear projection layer. For all other tasks, we retain the reconstruction head. We provide detailed descriptions of each task and MOMENT’s configuration in App. E.
@@ 4 Experimental Setup and Results
$\mathtt{MOMENT_{LP}}$ GPT4TS TimesNet N-BEATS ARIMA Theta ETS
Seasonal
Naive
Naive
Random
Walk
@@
For classification, we consider the unsupervised representation learning problem, where the goal is to learn representations of time series that are useful for downstream classification, without access to labeled data. As is common in prior work (Yue et al., 2022; Franceschi et al., 2019), the quality of representations is measured using the accuracy of a Support Vector Machine trained on them (App. E.2). For short-horizon forecasting, we consider the zero-shot setting introduced by Oreshkin et al. (2021). In particular, we fine-tune MOMENT on a source dataset using a forecasting head, and evaluate its performance on a target dataset without any fine-tuning (App E.1.2, Tab. 21).
Datasets. We use the same datasets as TimesNet for forecasting and imputation. However, for classification and anomaly detection, we conduct experiments on larger and systematically chosen subset of datasets from the UCR classification archive (Dau et al., 2018) and UCR anomaly archive (Wu & Keogh, 2023). Specifically, we run classification experiments on all 91 time series datasets with each time series shorter than 512 time steps (Tab.23). For anomaly detection, while choosing the subset of time series, we prioritized coverage over different domains and data sources represented in the UCR anomaly archive (Tab. 22). We also note that the UCR anomaly archive was proposed as an improvement over pre-existing anomaly detection datasets such as the SMD (Su et al., 2019), and SMAP (Hundman et al., 2018), many of which are also used in TimesNet. Our proposed experimental setup is summarized in Tab. 1 and detailed in App. E.
@@
[…]
@@
Wins/Losses 880.5/566.5 325.5/1121.5 326.0/1121.0 1220.0/227.0 1033.0/375.0 885.5/522.5 924.0/484.0 460.0/987.0 684.0/763.0 727.5/719.5 1031.5/415.5 533.0/914.0 696.5/750.5 1137.0/310.0 71.5/1375.5 593.0/854.0 712.5/734.5
Table 4: Classification accuracy of methods across 91 UCR datasets. Methods with mean and median accuracy higher than MOMENT are in bold. MOMENT without fine-tuning on individual datasets demonstrates promising accuracy. Complete results in Tab. 23.
@@ Acknowledgments
Discussions. We would like to express our sincerest gratitude to Barış Kurt, Andrey Kan, Laurent Callot, Gauthier Guinet, Jingchao Ni, and Jonas M. Kübler for insightful discussions regarding the problem setting and experimental design. Their unwavering support was instrumental in the development of MOMENT. We are also thankful to Laurent, Barış, Jingchao and Andrey for their constructive feedback on the writing of this manuscript. Additionally, we acknowledge the insightful exchanges with Yuyang (Bernie) Wang, Abdul Fatir Ansari, Ingo Guering, Xiyuan Zhang, and Anoop Deoras. Special thanks to Cherie Ho for suggesting a creative and befitting name for our model. Lastly, we would like to thank Cecilia Morales for her insightful comments, especially on the broader impacts of this work, and for helping us proofread this manuscript.
@@
All models were trained and evaluated on a computing cluster consisting of 128 AMD EPYC 7502 CPUs, 503 GB of RAM, and 8 NVIDIA RTX A6000 GPUs each with 49 GiB RAM. All MOMENT variants were trained on a single A6000 GPU (with any data or model parallelism). We have made MOMENT-large [^6] https://huggingface.co/AutonLab/MOMENT-1-large and the Time Series Pile [^7] https://huggingface.co/datasets/AutonLab/Timeseries-PILE publicly available on Huggingface. We are working on open-sourcing MOMENT-base and MOMENT-small, and our research code public. The latter is currently available anonymously at https://anonymous.4open.science/r/BETT-773F/README.md. We enlist an exhaustive list of hyper-parameters in App. E to aid reproducibility. We would like to emphasize that all datasets used in this study are publicly available.
@@
Ethical considerations and potential misuse. Despite MOMENT’s promising performance in limited-data settings, it is important to use its predictions with care, especially in high-stakes settings such as healthcare. Before MOMENT is used for high-stakes decision-making, we recommend fine-tuning and evaluating the model with task-specific in-domain data.
Appendix A Related Work
@@
Appendix B Contemporary Work
Feature MOMENT Moirai Lag-LLaMa Chronos TimesFM TimeGPT-1 TinyTimeMixers TOTEM
Citation Ours (Woo et al., 2024) (Rasul et al., 2023) (Ansari et al., 2024) (Das et al., 2023) (Garza & Mergenthaler-Canseco, 2023) (Ekambaram et al., 2024) (Talukder et al., 2024)
Base Architecture T5 encoder Encoder-only transformer LLaMa T5 (encoder-decoder) Decoder-only Transformer MLP-Mixer (Tolstikhin et al., 2021) VQ-VAE (Van Den Oord et al., 2017)
Sizes 40M, 125M, 385M 14M, 91M, 311M 200M 20M, 46M, 200M, 710M 17M, 70M, 200M ? $<$1M 0.3M
Open Source $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
@@ Appendix B Contemporary Work
Forecasting Forecasting Forecasting Forecasting Forecasting Forecasting
Forecasting, Classification,
Anomaly detection, Imputation
Data Total Observations 1.13B++ 27.65B [^9] These values are taken from Tables 8–17 of (Woo et al., 2024) and denote the # of observations 0.36B [^10] These values are taken from Table 4 of (Rasul et al., 2023) and denote the # of observations 84B 100B 100B 244M N/A
# Time Series 13M 4M 11K 11M ? ? N/A N/A
Public $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Public Release Upon acceptance Feb’2024 Feb’2024 Mar’2024 - - Apr’2024 Feb’2024
Tokenization Fixed-length patches Multi-scale patches Lag features Scaling & quantization Fixed-length patches ? Adaptive Patching VQVAE-based
@@
Forecast Cross-entropy loss Forecasting error (MSE) ? Forecasting error (MSE)
@@
Distributional Prediction In progress (post hoc) $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ (post hoc) $\times$ $\checkmark$
Interpretability $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
Any-variate Prediction $\checkmark$ (Channel independence) $\checkmark$ (Flattening) $\times$ $\times$ $\times$ $\checkmark$ (?) $\checkmark$ (Channel Independence + Mixing) $\checkmark$ (Flattening)
Context length 512 1000–5000 1024 512 512 ? 512 512
Scaling Model $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Data $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Impact of Initialization $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
@@
Appendix C Interesting directions for future work
@@
  • Study the impact of design choices such as the impact of the choice of the loss function (Huber, $L_{1}$, $L_{2}$), patch length (4,8)(4,8)( 4 , 8 ), and masking percentage (0.3,0.6)(0.3,0.6)( 0.3 , 0.6 ) on pre-training loss and time series modeling performance.
@@
Appendix D The Time Series Pile
@@
Long
horizon
forecasting
ETTm1, ETTm2 7 {96, 720} (33953, 11425, 11425) Electricity (15 mins)
@@
Short
horizon
forecasting
M4-Yearly 1 6 (16099, 2301, 4600) -
@@
Imputation
ETTm1, ETTm2 7 512 (33953, 11425, 11425) Electricity (15 mins)
@@
Classification
UWaveGestureLibraryX 1 315 (640, 256, 3582) Motion Gesture (8 classes)
@@
Anomaly
detection
(TSB-UAD)
1sddb40 1 - (24489, 9489, 3969) Beats
@@
Appendix E Experimental Setup and Results
@@ E.1 Forecasting
E.1.1 Long-Horizon Forecasting
@@
MOMENT sequence length : 512512512
patch length : 888
patch stride length : 888
initial learning rate : 0.00010.00010.0001
[…]
Time-LLM sequence length : 512512512
patch length : 161616
patch stride length : 888
initial learning rate : 0.0010.0010.001
dimension of feedforward layer : 204820482048
llm layers : 323232
number of heads : 888
N-BEATS sequence length : 512512512
@@
number of blocks per stack : 333
[…]
hidden layer units : 256256256
@@
E.1.2 Zero-shot Short-Horizon Forecasting
@@
$\mathtt{MOMENT_{LP}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
initial learning rate : 0.0020.0020.002
[…]
$\mathtt{MOMENT_{0}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
initial learning rate : 0.0010.0010.001
N-BEATS sequence length : 512512512
@@
number of blocks per stack : 333
[…]
hidden layer units : 256256256
GPT4TS forecast horizon : 0
gpt layers : 333
patch length : 111
patch stride length : 111
sequence length : 512512512
TimesNet sequence length : 512512512
model dimension : 323232
dimension of feedforward layer : 323232
[…]

### v3

@@ 1 Introduction
Figure 1: MOMENT can solve multiple time series analysis tasks well (App. F).
@@ 3.2 Model Architecture
Our transformer encoder retains the modifications proposed by Raffel et al. (2020) to the original Transformer (Vaswani et al., 2017). Specifically, we remove the additive bias from the Layer Norm (Ba et al., 2016), and place it before the residual skip connections (He et al., 2016), and use the relative positional embedding scheme (Shaw et al., 2018). Below we summarize the intuition behind some of our key design decisions.
@@
Additional absolute positional embeddings. In addition to relative positional embeddings, we add absolute sinusoidal positional embeddings (Vaswani et al., 2017) to each patch [^3] We initially included absolute positional embeddings by accident, but subsequent experiments showed that combining absolute and relative positional embeddings improve predictions..
@@
Reconstruction-based anomaly detection with window size $=512$
@@ 3.3 Pre-training using Masked Time series Modeling
We use the Adam optimizer with weight decay (Loshchilov & Hutter, 2019) with $\lambda=$ 0.05, $\beta_{1}=$ 0.9, $\beta_{2}=$ 0.999. We clip the gradient at 5.0, train models using a batch size of 2048, and use cosine learning rate schedule with initial and final learning rates of $1e^{-4}$ and $1e^{-5}$, respectively. We use gradient checkpointing (Radford et al., 2021) to improve training throughput and save memory, and train all models in a mixed precision setting, using float-32 for numerically unstable operations, e.g. layer normalization, and bfloat-16555https://cloud.google.com/tpu/docs/bfloat16, otherwise. We train all models for 2 epochs.
@@ 3.4 Fine-tuning on Downstream Tasks
MOMENT can be seamlessly used for multiple time series analysis tasks. In this work, we consider 5 practical time series analysis tasks as examples, namely: long- and short-horizon forecasting, classification, anomaly detection, and imputation. For forecasting tasks with horizon $H$, we replace the reconstruction head with a forecasting head, which first flattens all the $N$ $D$-dimensional patch embeddings into a $N\times D$ dimensional vector, and then projects it into a $H$-dimensional time series via a linear projection layer. For all other tasks, we retain the reconstruction head. We provide detailed descriptions of each task and MOMENT’s configuration in App. F.
@@ 4 Experimental Setup and Results
$\mathtt{MOMENT_{LP}}$ GPT4TS TimesNet N-BEATS ARIMA Theta ETS Seasonal Naive Naive Random Walk
@@
For classification, we consider the unsupervised representation learning problem, where the goal is to learn representations of time series that are useful for downstream classification, without access to labeled data. As is common in prior work (Yue et al., 2022; Franceschi et al., 2019), the quality of representations is measured using the accuracy of a Support Vector Machine trained on them (App. F.2). For short-horizon forecasting, we consider the zero-shot setting introduced by Oreshkin et al. (2021). In particular, we fine-tune MOMENT on a source dataset using a forecasting head, and evaluate its performance on a target dataset without any fine-tuning (App F.1.2, Tab. 21).
Datasets. We use the same datasets as TimesNet for forecasting and imputation. However, for classification and anomaly detection, we conduct experiments on larger and systematically chosen subset of datasets from the UCR classification archive (Dau et al., 2018) and UCR anomaly archive (Wu & Keogh, 2023). Specifically, we run classification experiments on all 91 time series datasets with each time series shorter than 512 time steps (Tab.23). For anomaly detection, while choosing the subset of time series, we prioritized coverage over different domains and data sources represented in the UCR anomaly archive (Tab. 22). We also note that the UCR anomaly archive was proposed as an improvement over pre-existing anomaly detection datasets such as the SMD (Su et al., 2019), and SMAP (Hundman et al., 2018), many of which are also used in TimesNet. Our proposed experimental setup is summarized in Tab. 1 and detailed in App. F.
@@
We do not perform hyper-parameter tuning. In all experiments that follow, unless mentioned otherwise, we fine-tune MOMENT-Large with a batch size of 64, and one cycle learning rate schedule with a peak learning rate between $55e-$ 5 and $31e-$ 3 (Smith & Topin, 2019). For baseline methods, we capture recommended settings from their papers and public repositories. We report all hyper-parameters settings for MOMENT and baselines in App. F.
@@
Wins 880 325 326 1220 1033 885 924 460 684 727 1031 533 696 1137 71 593 712
Losses 566 1121 1121 227 375 522 484 987 763 719 415 914 750 310 1375 854 734
Table 4: Classification accuracy of methods across 91 UCR datasets. Methods with mean and median accuracy higher than MOMENT are in bold. MOMENT without fine-tuning on individual datasets demonstrates promising accuracy. Complete results in Tab. 23. Results on multi-variate UEA datasets are in Tab. 24.
@@ Acknowledgments
Discussions. We would like to express our sincerest gratitude to Barış Kurt, Andrey Kan, Laurent Callot, Gauthier Guinet, Jingchao Ni, and Jonas M. Kübler for insightful discussions regarding the problem setting and experimental design. Their unwavering support was instrumental in the development of MOMENT. We are also thankful to Laurent, Barış, Jingchao and Andrey for their constructive feedback on the writing of this manuscript. Additionally, we acknowledge the insightful exchanges with Yuyang (Bernie) Wang, Abdul Fatir Ansari, Ingo Guering, Xiyuan Zhang, and Anoop Deoras. Special thanks to Cherie Ho for suggesting a creative and befitting name for our model. Lastly, we would like to thank Cecilia Morales for her insightful comments, especially on the broader impacts of this work, and for helping us proofread this manuscript. Finally, we would like to thank the anonymous ICML reviewers for their comments which helped improve our work.
@@
All models were trained and evaluated on a computing cluster consisting of 128 AMD EPYC 7502 CPUs, 503 GB of RAM, and 8 NVIDIA RTX A6000 GPUs each with 49 GiB RAM. All MOMENT variants were trained on a single A6000 GPU (with any data or model parallelism). We have made MOMENT-large [^7] https://huggingface.co/AutonLab/MOMENT-1-large and the Time Series Pile [^8] https://huggingface.co/datasets/AutonLab/Timeseries-PILE publicly available on Huggingface. We are working on open-sourcing MOMENT-base and MOMENT-small, and our research code public. The latter is currently available at https://github.com/moment-timeseries-foundation-model/moment-research. We enlist an exhaustive list of hyper-parameters in App. F to aid reproducibility. We would like to emphasize that all datasets used in this study are publicly available.
@@
Ethical considerations and potential misuse. Despite MOMENT’s promising performance in limited-data settings, it’s important to use its predictions with care, especially in high-stakes settings like healthcare. The reliability of MOMENT’s predictions hinges on the quality and diversity of the data it was trained on. Any dependence on unreliable or biased data could potentially result in skewed outputs, unfairly portraying or discriminating against certain groups or individuals. While MOMENT demonstrates the ability to capture changes time-series characteristics, e.g. trend, amplitude, frequency, it struggles to differentiate between vertically shifted time-series as it normalizes signals prior to modeling. Greater interpretability and explainability of MOMENT’s predictions are necessary to enable more informed decision-making processes. Prior to utilizing MOMENT in high-stakes decision-making, we recommend fine-tuning and evaluating the model with task-specific, in-domain data to address these ethical concerns and ensure equitable outcomes. Additionally, efforts should be made to promote equitable access to MOMENT and other AI technologies, particularly within healthcare systems where disparities in access may exacerbate existing inequities.
Appendix A Changelog
  • (Oct’2024) Added links to both research and inference benchmarks, notes on contemporary work in Tab. 9 (Timer (Liu et al., 2024) and TOTEM (Talukder et al., 2024)). We also note differences in positional embeddings (Sec. 3.2).
Appendix B Related Work
@@
Appendix C Contemporary Work
Feature MOMENT Moirai Lag-LLaMa Chronos Timer
Citation Ours (Woo et al., 2024) (Rasul et al., 2023) (Ansari et al., 2024) (Liu et al., 2024)
Base Architecture T5 encoder Encoder-only transformer LLaMa T5 (encoder-decoder) Decoder-only
Sizes 40M, 125M, 385M 14M, 91M, 311M 200M 20M, 46M, 200M, 710M 29M, 50M, 67M
Open Source $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ Appendix C Contemporary Work
Forecasting Forecasting Forecasting
Forecasting, Imputation,
Anomaly detection
Data Total Observations 1.13B++ 27.65B [^10] These values are taken from Tables 8–17 of (Woo et al., 2024) and denote the # of observations 0.36B [^11] These values are taken from Table 4 of (Rasul et al., 2023) and denote the # of observations 84B 28B
# Time Series 13M 4M 11K 11M 7.4M
Public $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Public Release May’24 Feb’2024 Feb’2024 Mar’2024 Jun’2024
Tokenization Fixed-length patches Multi-scale patches Lag features Scaling & quantization Single-series Sequence
@@
Forecast Cross-entropy loss Next token prediction
Distributional Prediction In progress (post hoc) $\checkmark$ $\checkmark$ $\checkmark$ $\times$
Interpretability $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Any-variate Prediction $\checkmark$ (Channel independence) $\checkmark$ (Flattening) $\times$ $\times$ $\checkmark$ (Uniform sub-sampling)
Context length 512 1000–5000 1024 512 1440
Scaling Model $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Data $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
Impact of Initialization $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Feature TimesFM TimeGPT-1 TinyTimeMixers TOTEM
Citation (Das et al., 2023) (Garza & Mergenthaler-Canseco, 2023) (Ekambaram et al., 2024) (Talukder et al., 2024)
Base Architecture Decoder-only Transformer MLP-Mixer (Tolstikhin et al., 2021) VQ-VAE (Van Den Oord et al., 2017)
Sizes 17M, 70M, 200M ? $<$1M 0.3M
Open Source $\times$ $\times$ $\checkmark$ $\checkmark$
Evaluation Tasks Forecasting Forecasting Forecasting
Forecasting, Classification,
Anomaly detection, Imputation
Data Total Observations 100B 100B 244M N/A
# Time Series ? ? N/A N/A
Public $\times$ $\times$ $\checkmark$ $\checkmark$
Public Release - - Apr’2024 Feb’2024
Tokenization Fixed-length patches ? Adaptive Patching VQVAE-based
Objective Forecasting error (MSE) ? Forecasting error (MSE)
@@
Distributional Prediction $\times$ $\checkmark$ (post hoc) $\times$ $\checkmark$
Interpretability $\checkmark$ $\times$ $\times$ $\times$
Any-variate Prediction $\times$ $\checkmark$ (?) $\checkmark$ (Channel Independence + Mixing) $\checkmark$ (Flattening)
Context length 512 ? 512 512
Scaling Model $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Data $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Impact of Initialization $\times$ $\times$ $\checkmark$ $\times$
@@
Appendix D Interesting directions for future work
@@
  • Study the impact of design choices such as the impact of the choice of the loss function (Huber, $L_{1}$, $L_{2}$), patch length $(4,8)$, and masking percentage $(0.3,0.6)$ on pre-training loss and time series modeling performance.
@@
Appendix E The Time Series Pile
@@
Long horizon forecasting (Informer) ETTm1, ETTm2 7 {96, 720} (33953, 11425, 11425) Electricity (15 mins)
@@
Short horizon forecasting (Monash) M4-Yearly 1 6 (16099, 2301, 4600) -
@@
Imputation (Informer) ETTm1, ETTm2 7 512 (33953, 11425, 11425) Electricity (15 mins)
@@
Classification (UCR) UWaveGestureLibraryX 1 315 (640, 256, 3582) Motion Gesture (8 classes)
@@
Anomaly detection (TSB-UAD) 1sddb40 1 - (24489, 9489, 3969) Beats
@@
Appendix F Experimental Setup and Results
@@ F.1 Forecasting
F.1.1 Long-Horizon Forecasting
@@
MOMENT sequence length : $512$
patch length : $8$
patch stride length : $8$
initial learning rate : $0.0001$
forecast horizon : $\{96,720\}$
Time-LLM sequence length : $512$
patch length : $16$
patch stride length : $8$
initial learning rate : $0.001$
dimension of feedforward layer : $2048$
llm layers : $32$
number of heads : $8$
N-BEATS sequence length : $512$
@@
number of blocks per stack : $3$
thetas dimensions : $\{4,8\}$
hidden layer units : $256$
@@
F.1.2 Zero-shot Short-Horizon Forecasting
@@
$\mathtt{MOMENT_{LP}}$ sequence length : $512$
patch length : $8$
patch stride length : $8$
initial learning rate : $0.002$
max epochs : $\{5,10\}$
$\mathtt{MOMENT_{0}}$ sequence length : $512$
patch length : $8$
patch stride length : $8$
initial learning rate : $0.001$
N-BEATS sequence length : $512$
@@
number of blocks per stack : $3$
thetas dimensions : $\{4,8\}$
hidden layer units : $256$
GPT4TS forecast horizon : 00
gpt layers : $3$
patch length : $1$
patch stride length : $1$
sequence length : $512$
TimesNet sequence length : $512$
model dimension : $32$
dimension of feedforward layer : $32$
top-$k$ : $5$